OpenStudy (cp9454):

A second ball is dropped from an elevator accelerating up with 1m/s2, 1 second after the first ball is dropped. how does relative velocity of two ball changes with time?

whitemonsterbunny17 (whitemonsterbunny17):

Sorry, but I have not learned Physics yet. Maybe @Abhisar can help. (:

OpenStudy (cp9454):

@LolWolf

OpenStudy (anonymous):

Increases linearly.

OpenStudy (cp9454):

relative velocity increases linearly?

OpenStudy (abhisar):

are u sure @LolWolf ? Can u xplain it why ?

OpenStudy (anonymous):

Wait, no, that's their distance. Clearly I'm not thinking today. Their relative velocity remains constant.

OpenStudy (abhisar):

\(\huge\checkmark\)

OpenStudy (abhisar):

U got that @cp9454 ?

OpenStudy (cp9454):

reason why relative velocity is constant.

OpenStudy (anonymous):

Reason (for the sake of completeness): \[ v(t)=at+v_0 \]And: \[ v'(t)=a(t+1) \]Then their relative velocity \[ v_r=v'(t)-v(t)=1a-v_0 \] Note that \(v_0<0\).

OpenStudy (abhisar):

Correct ! because at the moment the ball is dropped(separated from the lift) its accelaration becomes 9.8 or g

OpenStudy (abhisar):

This accelaration is same for both the balls.

OpenStudy (abhisar):

Now u can see by the relation @LolWolf provided that the relative velocity will remain constant !

OpenStudy (cp9454):

ok got it

OpenStudy (abhisar):

\(\huge\color{green}{ت}\)